# Post-hoc Popularity Bias Correction in GNN-based Collaborative Filtering

### v1

@@ Abstract.
collaborative filtering, graph neural networks, popularity bias
This is a preprint version.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; XXX XX–$X$, 2025; XXX, XXisbn: 978-1-4503-XXXX-X/XXXX/XXccs: Information systems Recommender systemscopyright: none
@@ 1. Introduction
Recommender systems are integral to modern online platforms, shaping user experiences through personalized recommendations across various domains, including e-commerce, entertainment, and social networking. By leveraging historical user–item interactions, they present users with relevant content from large item collections, a process commonly referred to as collaborative filtering (CF) (He et al., 2017; Su and Khoshgoftaar, 2009). Recently, Graph Neural Networks (GNNs) have emerged as a powerful method for CF by modeling recommendations as a user–item bipartite graph and propagating information to multi-hop neighborhoods via neighborhood aggregation (Yang et al., 2021). The aggregation mechanism enables GNN-based CF methods to learn higher-order interaction patterns, resulting in high-quality user and item representations and state-of-the-art performances (He et al., 2020; Gao et al., 2022b).
Recommender systems often rely on user interaction data, such as clicks, to infer preferences, since this data can easily be collected from interaction logs. However, such data is prone to biases (Ai et al., 2018; Joachims et al., 2017), with popularity bias being particularly severe (Zhou et al., 2023). Interaction data in recommender systems is often long-tailed, with a small set of items receiving most feedback while the majority remain rarely interacted with (Zhang et al., 2022b). CF models trained on such skewed data inherit popularity bias (Yao and Huang, 2017; Cañamares and Castells, 2018), over-recommending head items and under-recommending tail items. Through the feedback loop in recommender systems (Chaney et al., 2018), this bias can be amplified over time, reducing personalization and weakening long-term performance. Figure 1 illustrates the effect of data imbalance in interaction data. In the KuaiRec, Coat, and Yahoo! R3 datasets, items are sorted by interaction frequency, with the top $20\%$ categorized as head items and the remaining $80\%$ as tail items. User interactions are disproportionately concentrated on head items. Consequently, training models on such skewed data amplifies head-item recommendations, further marginalizing tail items in future recommendations.
@@
Several approaches have been proposed to mitigate popularity bias. A common strategy is to adjust the training loss to reduce the influence of popular items. Inverse propensity weighting (IPW) methods (Gruson et al., 2019; Joachims et al., 2017; Ai et al., 2018) reweight the loss by upweighting niche items and downweighting popular ones, while regularization-based methods (Boratto et al., 2021; Wasilewski and Hurley, 2016; Chen et al., 2020b) penalize correlations between predicted scores and item popularity. Causal graph-based methods (Ning et al., 2024; Wei et al., 2021; Bonner and Vasile, 2018; Wang et al., 2021) instead use causal structures to identify sources of popularity and apply causal inference to mitigate their effect on relevance estimation. However, all three of these loss-based approaches may not fully eliminate popularity bias in GNN-based CF (Zhou et al., 2023).
@@
To address this, recent works (Kim et al., 2022; Zhou et al., 2023) weight interactions during aggregation, akin to propensity scores, to downweight popular item interactions during aggregation. However, these methods assign weights in every training epoch—either using pre-computed scores (Kim et al., 2022) or dynamically estimated values (Zhou et al., 2023). Since model representations are unstable in the early stages, such weighting may distort the learning process. Moreover, both loss-based and aggregation-weight methods require retraining the backbone, making them impractical for deployed models (Chen et al., 2020a). Post-processing re-ranking methods (Abdollahpouri et al., 2019; Zhu et al., 2021) adjust predicted scores to diversity or popularity bias but cannot remove popularity effect embedded in node representations. More recently, a post-hoc method (Chen et al., 2024) for GNN-based CF estimates popularity effects using simple node degrees, without considering personalized preferences or interaction-level popularity effects. One might assume that debiasing can be achieved by simply removing some popular items from the recommendations to match the true underlying distribution. This distribution, however, is unknown in practice. Moreover, since popularity is not always harmful (Zhao et al., 2022), indiscriminately suppressing head items risks discarding truly relevant ones that reflect users’ true interests.
In this paper, we propose a Post-hoc Popularity Debiasing (PPD) method to mitigate popularity bias in GNN-based CF. Our method operates on pre-trained embeddings from an already trained model. We estimate the popularity of each user-item interaction by combining global preference with personalized preference derived from the embeddings. Based on these estimates, we identify a popularity direction vector for each node. We then project the node embedding onto this direction to obtain its component along it and update the embedding by removing this popularity component. This eliminates the popularity effect while preserving preference-driven signals, resulting in unbiased node representations for generating unbiased recommendations. Though this general idea is simple, this method provides a post-hoc solution that is very effective at reducing popularity bias without retraining the model.
@@ 2. Related Work
Recommender systems often exhibit popularity bias, causing popular items to dominate recommendations (Chen et al., 2022; Zhang et al., 2022b; Chen et al., 2020b; Rhee et al., 2022). Existing methods to address this include IPW, post-processing re-ranking, regularization, and causal inference methods.
@@
IPW-based methods mitigate popularity bias by down-weighting frequent items, assigning each interaction a weight inversely proportional to the item’s popularity (Ai et al., 2018; Joachims et al., 2017). To reduce the high variance from such reweighting, subsequent works (Gruson et al., 2019; Bottou et al., 2013) use normalization or smoothing penalties. However, the effectiveness of IPW depends on accurate propensity score estimation (Ovaisi et al., 2021). IPW is also widely used in learning-to-rank systems to correct biases such as position and selection bias (Joachims et al., 2017; Oosterhuis and de Rijke, 2020; Wang et al., 2016; Ai et al., 2018; Luo et al., 2024, 2023). Regularization-based methods incorporate constraints during training to reduce popularity effects. LapDQ (Wasilewski and Hurley, 2016) promotes diversity via item-distance regularization, INRS (Kamishima et al., 2014) applies mean-matching to favor less popular items, ESAM (Chen et al., 2020b) transfers knowledge from popular to unpopular items to alleviate sparsity, and Sam-Reg (Boratto et al., 2021) reduces the correlation between relevance and popularity for fairer treatment. Post-processing re-ranking methods operate after recommendation generation, leaving user and item representations intact. They address goals such as preserving historical preference distributions (calibration) (Steck, 2018), enhancing diversity (Abdollahpouri et al., 2019), and correcting popularity bias by adjusting predicted scores (Zhu et al., 2021). Causal inference has been used to separate user interests from popularity effects (Bonner and Vasile, 2018; Wang et al., 2021; Wei et al., 2021; Ning et al., 2024; He et al., 2022; Zhao et al., 2022). CausE (Bonner and Vasile, 2018) uses causal graphs for unbiased learning, MPCI (He et al., 2022) blocks paths from popularity to predictions, MACR (Wei et al., 2021) removes direct popularity effects via multi-task counterfactual learning, and PPAC (Ning et al., 2024) jointly models global and personal popularity within user neighborhoods using counterfactual inference.
@@
Deep learning-based recommender systems rely on high-quality representations to predict user–item interactions (Zha et al., 2022). GNNs achieve this by aggregating information from neighboring nodes (Chen et al., 2022, 2020c; Dong et al., 2023; Huang et al., 2022; Zhang et al., 2023a, 2022a), with propagation layers capturing higher-order connectivities. They are state-of-the-art for CF due to this ability. LightGCN (He et al., 2020) simplifies GNNs by removing non-linear transformations and activations and is widely used in GNN-based CF. NIA-GCN (Sun et al., 2020) models pairwise neighbor relationships, while UltraGCN (Mao et al., 2021) addresses over-smoothing by approximating infinite-layer limits. With the rise of self-supervised learning, methods like SGL (Wu et al., 2021) and NCL (Lin et al., 2022) enhance robustness by combining contrastive objectives with structural and semantic signals, while DirectAU (Wang et al., 2022) emphasizes alignment and uniformity in representations. SimGCL (Yu et al., 2022) and XSimGCL (Yu et al., 2023) further show that contrastive learning improves recommendations by promoting uniform user–item embeddings and propose a simple noise-based perturbation strategy. However, these approaches overlook popularity amplification in GNNs. APDA (Zhou et al., 2023) addresses this by reweighting interactions using inverse popularity scores during aggregation, and NAVIP (Kim et al., 2022) applies IPW during aggregation based on interaction counts.
Recent studies have explored out-of-distribution (OOD) generalization in GNN-based recommendation. DRO (Zhao et al., 2025b) reduces the impact of noisy samples, CausalDiffRec (Zhao et al., 2025a) uses causal diffusion to remove environmental confounders, and AdvInfoNCE (Zhang et al., 2023b) employs a hardness-aware adversarial contrastive loss to improve generalization. Despite their effectiveness, these methods do not explicitly address popularity bias. DAP (Chen et al., 2024), a post-hoc popularity debiasing method, estimates popularity using both clustering of similar nodes and information from high- and low-degree neighbors, and removes popularity from node embeddings.
@@ 3.1. Preliminaries
For example, the $f_{\text{readout}}(\cdot)$ function in LightGCN is a mean-pooling operation that combines information from all layers to form the final representation of each node: $\mathbf{e}_{u}=\frac{1}{L+1}\sum_{l=0}^{L}\mathbf{e}_{u}^{(l)},\mathbf{e}_{i}=\frac{1}{L+1}\sum_{l=0}^{L}\mathbf{e}_{i}^{(l)}.$ The predicted score between a user $u$ and an item $i$ is computed as the inner product of user and item final representations, i.e., ${\hat{y}}_{ui}=\mathbf{e}_{u}^{\top}\mathbf{e}_{i}$. Using these scores, the top-$k$ items that user $u$ has not interacted with can be recommended, which are most likely to be clicked by user $u$.
In general, the most commonly used objective function in GNN-based CF (e.g., LightGCN) is the pairwise Bayesian Personalized Ranking (BPR) loss (Rendle et al., 2012), which maximizes the difference between positive and negative item scores with an $L_{2}$ regularization term to prevent overfitting:
@@ 4.1. Popularity Estimation
User-item interactions can arise from two main factors, namely the global preference of the item and the user’s personalized preference for it. To quantify the contribution of popularity to each interaction, we define three measures: global preference, personalized preference, and popularity score. Global preference reflects the fact that some items receive disproportionately high recommendations across the user base. To capture this effect in the embedding space, we measure how similar an item is, on average, to all users. Items with higher similarity tend to obtain higher prediction scores and are more likely to be recommended, reflecting the extent to which interactions are driven by broader popularity.
@@
Personalized preference reflects how well an item matches the behavior of a specific user. If an interacted item is similar to a user’s historical interactions, it is likely to reflect the user’s personalized preferences. However, since popular items may also appear in a user’s history, we adjust their influence when computing personalized preferences.
@@
The popularity score $b_{ui}$ for each interaction estimates the extent to which an interaction between user $u$ and item $i$ is influenced by popularity rather than user preference. The reason for subtracting personalized preference is that some part of an item’s global preference reflects true user preferences. By removing this component, the popularity score captures only the residual effect of popularity, isolating the part of the interaction that is driven purely by popularity. A high score indicates that the item is popular but not well-aligned with the user’s interests, suggesting a popularity-driven interaction. A low score implies alignment with the user’s preferences, reflecting a more true preference. To make the two scores $p_{i}$ and $r_{ui}$ comparable, we can normalize them before computing $b_{ui}$. We compute the popularity score for each interaction using the stable user and item embeddings learned by the pre-trained model. This yields a more accurate score, as the embeddings already capture both popularity and meaningful user interests. This is an advantage of our post-hoc approach over methods (Zhou et al., 2023; Kim et al., 2022) that estimate popularity during training.
@@ 4.2. Post-hoc Popularity Bias Correction via Projection
The popularity direction vector captures how popularity pulls the node embeddings away from preferences. This vector represents the direction from preference to popularity in embedding space. The preference centroid coefficient $\phi\in[0,1]$ controls the contribution of the preference centroid in the popularity direction. In datasets where popularity bias is severe due to heavily skewed data, preference signals may contain unreliable or less meaningful information. In such cases, using a larger $\phi$ may yield a less accurate estimate of the popularity direction, whereas a smaller $\phi$ can provide a more reliable estimation. In datasets with more diverse interactions and less popularity, preference signals can be more reliable, and a larger $\phi$ enables a more accurate estimation of the popularity direction.
We leverage the vector projection formula (Perwass, 2009) to decompose a vector along the direction of another. By projecting one vector onto another, we can obtain the component of the first vector that lies along the direction of the second (Perwass, 2009). Accordingly, we project the node embedding onto the popularity direction vector to get the component of an embedding aligned with that direction. By subtracting this component from the node embedding, we remove the popularity component while preserving user preferences, and thus can obtain unbiased embeddings for users and items free of popularity. The final debiased embeddings for user $u$ at layer 0 can be written as:
@@
These debiased embeddings are used as input to the propagation layers of the model. The information propagation for layers $l\geq 1$ can be computed using equation (1), where $\mathbf{e}_{u}^{(0)}:=\tilde{\mathbf{e}}_{u}^{(0)}$ and $\mathbf{e}_{i}^{(0)}:=\tilde{\mathbf{e}}_{i}^{(0)}$ are the debiased embeddings at layer 0. By debiasing the 0th-layer embeddings and propagating them in higher layers, the model propagates unbiased node information to higher layers. Thus, we remove popularity at the 0-th layer to allow unbiased information to flow through higher layers. The PPD algorithm and its complexity analysis are given in Subsections A.1 and A.2 of the Appendix.
Finally, the final embedding of a node is obtained by combining information from all layers using equation (2). The predicted relevance score between user $u$ and item $i$ is the inner product of their final embeddings: $\hat{y}_{ui}=\mathbf{e}_{u}^{\top}\mathbf{e}_{i}$. Based on these scores, the model recommends the top-$k$ items not yet interacted with by user $u$ that are most likely to reflect their interests. Our approach is post-hoc, applying bias correction to already learned node representations. Thus, it can correct bias in a deployed model without retraining, updating only the embeddings of a trained model to mitigate popularity bias.

### v2

@@ Abstract.
collaborative filtering; graph neural networks; popularity bias
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792670isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Recommender systems
Resource Availability:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.18371801.
@@ 1. Introduction
Recommender systems are integral to modern online platforms, shaping user experiences through personalized recommendations across various domains, including e-commerce, entertainment, and social networking. By leveraging historical user–item interactions, they present users with relevant content from large item collections, a process commonly referred to as collaborative filtering (CF) (He et al., 2017; Su and Khoshgoftaar, 2009). Recently, Graph Neural Networks (GNNs) have emerged as a powerful method for CF by modeling recommendations as a user–item bipartite graph and propagating information to multi-hop neighborhoods via neighborhood aggregation (Yang et al., 2021). The aggregation mechanism enables GNN-based CF methods to learn higher-order interaction patterns, resulting in high-quality user and item representations and state-of-the-art performances (He et al., 2020; Gao et al., 2022a).
Recommender systems often rely on user interaction data, such as clicks, to infer preferences, since this data can easily be collected from interaction logs. However, such data is prone to biases (Ai et al., 2018; Joachims et al., 2017), with popularity bias being particularly severe (Zhou et al., 2023). Interaction data in recommender systems is often long-tailed, with a small set of items receiving most feedback while the majority remain rarely interacted with (Zhang et al., 2022a). CF models trained on such skewed data inherit popularity bias (Yao and Huang, 2017; Cañamares and Castells, 2018), over-recommending head items and under-recommending tail items. Through the feedback loop in recommender systems (Chaney et al., 2018), this bias can be amplified over time, reducing personalization and weakening long-term performance. Figure 1 illustrates the effect of data imbalance in interaction data. In the KuaiRec, Coat, and Yahoo! R3 datasets, items are sorted by interaction frequency, with the top $20\%$ categorized as head items and the remaining $80\%$ as tail items. User interactions are disproportionately concentrated on head items.
@@
Consequently, training models on such skewed data amplifies head-item recommendations, further marginalizing tail items in future recommendations.
Several approaches have been proposed to mitigate popularity bias. A common strategy is to adjust the training loss to reduce the influence of popular items. Inverse propensity weighting (IPW) methods (Gruson et al., 2019; Joachims et al., 2017; Ai et al., 2018) reweight the loss by upweighting unpopular items and downweighting popular ones, while regularization-based methods (Boratto et al., 2021; Wasilewski and Hurley, 2016; Chen et al., 2020c) penalize correlations between predicted scores and item popularity. Causal graph-based methods (Ning et al., 2024; Wei et al., 2021; Bonner and Vasile, 2018; Wang et al., 2021) instead use causal structures to identify sources of popularity and apply causal inference techniques to mitigate their effect on relevance estimation. However, all three of these loss-based approaches may not fully eliminate popularity bias in GNN-based CF (Zhou et al., 2023).
@@
To address this, recent works (Kim et al., 2022; Zhou et al., 2023) weight interactions during aggregation, akin to propensity scores, to downweight popular item interactions during aggregation. However, these methods assign weights in every training epoch—either using pre-computed scores (Kim et al., 2022) or dynamically estimated values (Zhou et al., 2023). Since model representations are unstable in the early stages, the resulting embeddings are under-trained and noisy, which leads to unreliable popularity estimates that can distort and mislead the learning process.
Both loss-based and aggregation-weight methods require retraining the backbone, making them impractical for deployed models (Chen et al., 2020a). Post-processing re-ranking methods (Abdollahpouri et al., 2019; Zhu et al., 2021) adjust predicted scores to diversity or popularity bias but cannot remove popularity effect embedded in node representations. More recently, a post-hoc method (Chen et al., 2024) for GNN-based CF estimates popularity effects using simple node degrees, without considering personalized preferences or interaction-level popularity effects. One might assume that debiasing can be achieved by simply removing some popular items from the recommendations to match the true underlying distribution. This distribution, however, is unknown in practice. Moreover, since popularity is not always harmful (Zhao et al., 2022), indiscriminately suppressing head items risks discarding truly relevant items of user interest.
In this paper, we propose a Post-hoc Popularity Debiasing (PPD) method to mitigate popularity bias in GNN-based CF. Our method operates on pre-trained embeddings from an already trained model. We estimate the popularity of each user-item interaction by combining global preference with personalized preference derived from the embeddings. Based on these estimates, we identify a popularity direction vector for each node. We then project the node embedding onto this direction to obtain its component along the direction and update the embedding by removing this popularity component. This eliminates the popularity effect while preserving preference-driven signals, resulting in unbiased node representations for generating unbiased recommendations. Though this general idea is simple, this method provides a post-hoc solution that is very effective at reducing popularity bias without retraining the model.
@@ 2. Related Work
Recommender systems often exhibit popularity bias, causing popular items to dominate recommendations (Chen et al., 2022; Zhang et al., 2022a; Chen et al., 2020c; Rhee et al., 2022). Existing methods to address this include IPW, post-processing re-ranking, regularization, and causal inference methods.
@@
IPW-based methods mitigate popularity bias by down-weighting frequent items, assigning each interaction a weight inversely proportional to the item’s popularity (Ai et al., 2018; Joachims et al., 2017). To reduce the high variance of IPW, subsequent works (Gruson et al., 2019; Bottou et al., 2013) use normalization or smoothing penalties. However, the effectiveness of IPW depends on accurate propensity score estimation (Ovaisi et al., 2021). IPW is also widely used in learning-to-rank systems to correct biases such as position and selection bias (Joachims et al., 2017; Oosterhuis and de Rijke, 2020; Wang et al., 2016; Ai et al., 2018; Luo et al., 2024, 2023). Regularization-based methods incorporate constraints during training to reduce popularity. LapDQ (Wasilewski and Hurley, 2016) promotes diversity via item-distance regularization, INRS (Kamishima et al., 2014) applies mean-matching to favor less popular items, ESAM (Chen et al., 2020c) transfers knowledge from popular to unpopular items to alleviate sparsity, and Sam-Reg (Boratto et al., 2021) reduces the correlation between relevance and popularity. Post-processing re-ranking methods operate after recommendation generation, leaving user and item representations intact. They address goals such as preserving historical preference distributions (calibration) (Steck, 2018), enhancing diversity (Abdollahpouri et al., 2019), and correcting popularity bias in dynamic recommender systems by adjusting predicted scores (Zhu et al., 2021). Causal inference has been used to separate user interests from popularity effects (Bonner and Vasile, 2018; Wang et al., 2021; Wei et al., 2021; Ning et al., 2024; He et al., 2022; Zhao et al., 2022). CausE (Bonner and Vasile, 2018) uses causal graphs for unbiased learning, MPCI (He et al., 2022) blocks paths from popularity to predictions, MACR (Wei et al., 2021) removes direct popularity effects via multi-task counterfactual learning, and PPAC (Ning et al., 2024) jointly models global and personal popularity within user neighborhoods using counterfactual inference.
@@
Deep learning-based recommender systems rely on high-quality representations to predict user–item interactions (Zha et al., 2022). GNNs achieve this by aggregating information from neighboring nodes (Chen et al., 2022, 2020b; Dong et al., 2023; Huang et al., 2022; Zhang et al., 2023b, 2022b), with propagation layers capturing higher-order connectivities. They are state-of-the-art for CF due to this ability. LightGCN (He et al., 2020) simplifies GNNs by removing non-linear transformations and activations and is widely used in GNN-based CF. NIA-GCN (Sun et al., 2020) models pairwise neighbor relationships, while UltraGCN (Mao et al., 2021) addresses over-smoothing by approximating infinite-layer limits. With the rise of self-supervised learning, methods like SGL (Wu et al., 2021) and NCL (Lin et al., 2022) enhance robustness by combining contrastive objectives with structural and semantic signals, while DirectAU (Wang et al., 2022) emphasizes alignment and uniformity in representations. SimGCL (Yu et al., 2022) and XSimGCL (Yu et al., 2023) further show that contrastive learning improves recommendations by promoting uniform user–item embeddings. However, these approaches overlook popularity amplification in GNNs. APDA (Zhou et al., 2023) addresses this by reweighting interactions using inverse popularity scores during aggregation, and NAVIP (Kim et al., 2022) applies IPW during aggregation based on interaction counts. Apart from these, Zhang et al. (Zhang et al., 2023c) propose a Cross Decoupling Network (CDN) that decouples memorization and user distributions to mitigate long-tail bias in content-based recommender systems. Recent studies have explored out-of-distribution (OOD) generalization in GNN-based recommendation. DRO (Zhao et al., 2025b) mitigates the effect of noisy samples, CausalDiffRec (Zhao et al., 2025a) removes environmental confounders via causal diffusion, and AdvInfoNCE (Zhang et al., 2023a) improves generalization using a hardness-aware adversarial contrastive loss. Despite their effectiveness, these methods do not explicitly address popularity bias. DAP (Chen et al., 2024) is a post-hoc debiasing approach that estimates popularity via clustering similar nodes and leveraging signals from high- and low-degree neighbors, and then removes popularity information from node embeddings.
@@ 3.1. Preliminaries
For example, the $f_{\text{readout}}(\cdot)$ function in LightGCN is a mean-pooling operation that combines information from all layers to form the final representation of each node: $\mathbf{e}_{u}=\frac{1}{L+1}\sum_{l=0}^{L}\mathbf{e}_{u}^{(l)},\mathbf{e}_{i}=\frac{1}{L+1}\sum_{l=0}^{L}\mathbf{e}_{i}^{(l)}.$ The predicted score between a user $u$ and an item $i$ is computed as the inner product of user and item final representations, i.e., ${\hat{y}}_{ui}=\mathbf{e}_{u}^{\top}\mathbf{e}_{i}$. Using these scores, the top-$k$ items that user $u$ has not interacted with can be recommended.
The most commonly used objective function in GNN-based CF (e.g., LightGCN) is the pairwise Bayesian Personalized Ranking (BPR) loss (Rendle et al., 2012), which maximizes the difference between positive and negative item scores with an $L_{2}$ regularization term:
@@ 4.1. Popularity estimation
User-item interactions can arise from two main factors, namely the global preference of the item and the user’s personalized preference for it. To quantify the contribution of popularity to each interaction, we define three measures: global preference, personalized preference, and popularity score.
4.1.1. Global preference
Global preference reflects the fact that some items receive disproportionately high recommendations across the user base. To capture this effect in the embedding space, we measure how similar an item is, on average, to all users. Items with higher similarity with users tend to obtain higher prediction scores and are more likely to be recommended, reflecting the extent to which interactions are driven by broader popularity.
@@ 4.1.1. Global preference
4.1.2. Personalized preference
Personalized preference reflects how well an item matches the historical behavior of a user. If an interacted item is similar to a user’s historical interactions, it likely reflects the user’s personalized preferences. However, since popular items may appear in a user’s history, we adjust their influence when computing personalized preferences.
@@ 4.1.3. Popularity score
The popularity score $b_{ui}$ for each interaction estimates the extent to which an interaction between user $u$ and item $i$ is influenced by popularity rather than user preference. The reason for subtracting personalized preference is that some part of an item’s global preference reflects true user preferences. Considering global preference ($p_{i}$) directly to address popularity bias can hurt recommendations (Zhao et al., 2022), as some popular items can truly be relevant. By removing personalized preference component, the popularity score captures only the residual effect of popularity, isolating the part of the interaction that is driven purely by popularity. Without this, preference-aligned popular items can incorrectly be treated as pure popularity. A high score indicates that the item is popular but not well-aligned with the user’s interests, suggesting a popularity-driven interaction. A low score implies alignment with the user’s preferences, reflecting a more true preference. To make the two scores $p_{i}$ and $r_{ui}$ comparable, we can normalize them before computing $b_{ui}$. We compute the popularity score for each interaction using the stable user and item embeddings learned by the pre-trained model. This yields a more accurate score, as the embeddings already capture both popularity and meaningful user interests. This is an advantage of our approach over methods (Zhou et al., 2023; Kim et al., 2022) that estimate popularity during training.
@@ 4.2. Post-hoc popularity bias correction via projection
The popularity direction vector captures how popularity pulls the node embeddings away from preferences. This vector represents the direction from preference to popularity in embedding space. The preference centroid coefficient $\phi\in[0,1]$ controls the contribution of the preference centroid in the popularity direction and adjusts for potential noise in the preference centroid. In datasets where popularity bias is severe due to heavily skewed data, preference signals may contain unreliable or noisy information. In such cases, using a larger $\phi$ may yield a less accurate estimate of the popularity direction, whereas a smaller $\phi$ can provide a more reliable estimation. In datasets with more diverse interactions and less popularity, preference signals can be more reliable, and a larger $\phi$ enables a more accurate estimate of the popularity direction.
We leverage the vector projection formula (Perwass, 2009) to decompose a vector along the direction of another. By projecting one vector onto another, we can obtain the component of the first vector that lies along the direction of the second (Perwass, 2009). We project the node embedding onto the popularity direction vector to get the component of an embedding aligned with that direction. By subtracting this component from the node embedding, we remove the popularity component while preserving user preferences, and thus can obtain unbiased embeddings for users and items free from popularity. The final debiased embeddings for user $u$ at layer 0 can be written as:
@@
These debiased embeddings are used as input to the propagation layers of the model. The information propagation for layers $l\geq 1$ can be computed using equation (1), where $\mathbf{e}_{u}^{(0)}:=\tilde{\mathbf{e}}_{u}^{(0)}$ and $\mathbf{e}_{i}^{(0)}:=\tilde{\mathbf{e}}_{i}^{(0)}$ are the debiased embeddings at layer 0. By debiasing the 0th-layer embeddings and propagating them in higher layers, the model propagates unbiased node information to higher layers. Thus, we remove popularity at the 0-th layer to allow unbiased information to flow through higher layers. The full PPD algorithm is provided in Subsection A.1 of the Appendix.
Finally, the final embedding of a node is obtained by combining information from all layers using equation (2). The predicted relevance score between user $u$ and item $i$ is the inner product of their final embeddings: $\hat{y}_{ui}=\mathbf{e}_{u}^{\top}\mathbf{e}_{i}$. Based on these scores, the model can recommend the top-$k$ items that user $u$ has not yet interacted with and that are most likely to reflect the user’s interests. Our approach is post-hoc, applying bias correction to already learned node representations. Thus, it can correct bias in a deployed model without retraining, updating only the embeddings of a trained model to mitigate popularity bias.
Note that PPD requires only the embeddings from a pre-trained model and can therefore be used as a general post-hoc debiasing strategy for other latent-factor models (e.g., matrix factorization) using the same procedure, except propagating information from unbiased embeddings. We focus on GNN-based CF in our current study because GNNs are state-of-the-art CF models but tend to severely amplify popularity bias through propagation (Zhou et al., 2023), making them an important use case. Exploring post-hoc methods for other latent-factor models can be a good future research direction. 00footnotetext: This work was supervised by Ren Wang and Elena Zheleva.